Brenda's school is selling tickets to a spring musical. On the first day of ticket sales the school sold 3 senior citizen tickets and 9 child tickets for a total of $75.00. The school took in $67.00 on the second day by selling 8 senior citizen tickets and 5 child tickets. What is the price each of one senior citizen ticket and one child ticket?
Answer by mananth(16946)

let senior citizen ticket cost $x
child ticket cost $y
..
First day
3x+9y=75
/3
x+3y=25..................1
..
Second Day
8x+5y=67.................2
multiply equation 1 by -8
-8x-24y=-200
add it to equation 2
-24y+5y=-200+67
-19y=-133
/-19
y= $7 cost of child ticket
plug value of y in equation1
x+21=25
x=$4 cost of child ticket
...
m.ananth@hotmail.ca